OpenStudy (anonymous):

Solve the following quadratic equation by completing the square: x2 + 4x – 1 = 0

sam (.sam.):

x2 + 4x – 1 = 0 (x+2)^2-4-1=0 (x+2)^2-5=0 add both sides 5 (x+2)^2=5 square root both sides \[x+2= \pm 5\] solve

OpenStudy (anonymous):

^ x + 2 = +/-sqrt(5)

OpenStudy (anonymous):